OpenStudy (anonymous):

The other day we were given a practice AB calculus exam (since I am in BC, we have already covered all of AB), and I had a quick question from it. I probably knew the answer to this, but forgot it, so if someone could give me a refresher by answering this question, that would be much appreciated.

OpenStudy (anonymous):

Why do the absolute values on ln\[\left[ x^2-1 \right]\] go away when you derive it?

OpenStudy (anonymous):

\[\ln \left[ x^2-1 \right]\]

OpenStudy (anonymous):

This was a multiple choice question, and they just wanted me to derive f(x). The answer is 2x/(x^2-1)

OpenStudy (anonymous):

the derivative of \[\ln(f(x))\] is, by the chain rule \[\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

Sorry, I was having wifi issues. I get it now! Thanks!!

OpenStudy (anonymous):

So just completely ignore the absolute values?